OpenStudy (anonymous):

2+3=8, 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8=?? Solve it?

OpenStudy (anonymous):

will give medal if solved

OpenStudy (automaticloveletter):

2+3=2*[3+(2-1)]=8 3+7=3*[7+(3-1)]=27 4+5=4*[5+(4-1)]=32 5+8=5*[8+(5-1)]=60 6+7=6*[7+(6-1)]=72 7+8=7*[8+(7-1)]=98 x+y=x[y+(x-1)]=x^2+xy-x

OpenStudy (automaticloveletter):

Thanks so much! all you have to do is add them and subtract by one. And the number you get, multiply it by the first number in the equation. :]